OpenStudy (anonymous):

Solve the following system of equations using elimination: 3x + 5y = 12 2x + 8y = 22 A. (-1, 3) B. (1, -3) C. (1, 3) D. (3, -1)

OpenStudy (anonymous):

you want to multiply either equation by a constant that will cancel out one variable from either equation when you subtract the equations, solve for the variable you have left and then plug that back into either equation to solve for the remaining variable

OpenStudy (anonymous):

I don't understand can you show me steps you don't have to give me the answer.

OpenStudy (anonymous):

in your case you want to multiply 3x+5y=12 by 2 on both sides of the equation and 2x+8y=22 by -3 on both sides.. you should end up with two new equations that look like 6x+10y=24 -6x-24y=-66 from here you can cancel out the x term by adding the equations and solving for y

OpenStudy (anonymous):

are you ok from there?

OpenStudy (anonymous):

I don't know...

OpenStudy (anonymous):

when you combine the two equations you get -14y=-42 from there you just solve for y by dividing by -14 on both sides y=3 now plug y into either of the equations to solve for x 3x+5y=12 .... 3x+5(3)=12 3x+15=12 3x=-3 divide by 3 both sides x=-1 so your answer is (-1,3)